Solve the following non-homogeneous linear ODE of first order

dy/dx + 3y/x =6x^2


1
Expert's answer
2021-04-15T07:21:36-0400

Solution

Let’s find solution of the homogeneous equation dy0/dx + 3y0/x = 0 in the form y0 = Cxk. Substitution into this equation gives

Ckxk-1 + 3Cxk-1 = 0  => k+3 = 0 => k = -3. = y0 = Cx-3.

Partial solution of nonhomogeneous equation let’s find as the expression y1 = Ax3. Substitution into equation gives

3Ax2 + 3Ax2 = 6x2 =>  A = 1 => y1 = x3.

Therefor the solution of the given equation is

y = y0+y1 = Cx-3+ x3

Answer

y = Cx-3+ x3